Camilla has three times as many points as Lynn, Lynn has five more points that Kim. Camilla, Lynn and Kim combined have twice as many points as Jasmine. If Jasmine has 25 points, how many points does each of the other three girls have?
Answer by stanbon(75887)

Camilla has three times as many points as Lynn,
Lynn has five more points that Kim.
Camilla, Lynn and Kim combined have twice as many points as Jasmine.
If Jasmine has 25 points, how many points does each of the other three girls have?
------------------
C = 3L
L = K + 5
C + L + K = 2J
J = 25
-------------------
J = 25
C + L + K = 50
0 + L - K = 5
C -3L + 0 = 0
-------------------
Solve the system of three equations with the three variables C,L, and K
I used the Matrix function on a TI-84 and got:
C = 33
L = 11
K = 6
------------------
Cheers,
Stan H.